OpenStudy (anonymous):

Find the product of the complex number and its conjugate. 1 + 2i

OpenStudy (anonymous):

(a+bi)(a-bi) = a^2 + b^2

OpenStudy (anonymous):

let z = 1 + 2i therefore it conjugate is given as conjugate z= 1-2i z*conjugate z = (1+2i)(1-2i) = 1^2 - (2i)^2 = 1- 4i^2= 1-4(-1) =1+4 = 5 ( since i^2 =-1)

OpenStudy (anonymous):

et z = 1 + 2i therefore it conjugate is given as conjugate z= 1-2i \[z*conjugate z = (1+2i)(1-2i) = 1^2 - (2i)^2 = 1- 4i^2= 1-4(-1) =1+4 = 5\] ( since \[i^2 =-1\])